Integrate x^2+X+5/(x^2+4)(X+1) dx
1
Expert's answer
2020-04-20T11:21:40-0400

"\u222b(x^2+x+5)\/((x^2+4)(x+1)) dx=\u2297"

"(x^2+x+5)\/((x^2+4)(x+1))=(Ax+B)\/(x^2+4)+C\/(x+1)" ;

"(Ax+B)(x+1)+C(x^2+4)=x^2+x+5;"

"x^2:A+C=1;"

"x:A+B=1;"

"x^0:B+4C=5;"

"C=1-A;B=1-A;1-A+4(1-A)=5;-5A=0;"

"A=0;B=1;C=1;"

"\u2297=\u222bdx\/(x^2+4)+ \u222bdx\/(x+1)= (1\/2) arctan (x\/2)+ln\u2061|x+1|+C."